the standard concentration of nitroglycerin is 25 mg/250mL. a person receives 250 mcg/min as an IV infusion. How many mL of the nitroglycerin solution has the patient received in 1.5 hours
Answer #1

person receives 250 mcg/min

time = 1.5 hr = 1.5 * 60 = 90 min.

thus

amount of IV infusion receives = 90 min * 250 mcg/min = 22500 mcg = 22500 / 1000 = 22.5 mg

and

mL of the nitroglycerin solution = 250 ml * 22.5 mg / 25 mg = 225 ml (answer)